# Limitations on Safe, Trusted, Artificial General Intelligence

### v1

@@ Abstract
Safety, trust and Artificial General Intelligence (AGI) are aspirational goals in artificial intelligence (AI) systems, and there are several informal interpretations of these notions. In this paper, we propose strict, mathematical definitions of safety, trust, and AGI, and demonstrate a fundamental incompatibility between them. We define safety of a system as the property that it never makes any false claims, trust as the assumption that the system is safe, and AGI as the property of an AI system always matching or exceeding human capability. Our core finding is that—for our formal definitions of these notions—a safe and trusted AI system cannot be an AGI system: for such a safe, trusted system there are task instances which are easily and provably solvable by a human but not by the system. We note that we consider strict mathematical definitions of safety and trust, and it is possible for real-world deployments to instead rely on alternate, practical interpretations of these notions. We show our results for program verification, planning, and graph reachability. Our proofs draw parallels to Gödel’s incompleteness theorems and Turing’s proof of the undecidability of the halting problem, and can be regarded as interpretations of Gödel’s and Turing’s results.
@@ 1 Introduction
Rapid advancements in artificial intelligence (AI) have intensified focus on Artificial General Intelligence (AGI) — loosely understood to be systems capable of human-level cognitive function across diverse tasks (Morris et al., 2024; Feng et al., 2024). AGI systems have the potential for vast societal benefits through transformative impacts on nearly every aspect of society, including healthcare (Singhal et al., 2025), scientific research (Wang et al., 2023), education (Wang et al., 2024), sustainability (Rolnick et al., 2022), and economic growth (Chui et al., 2023). At the same time, development of such powerful systems necessitates a foundational emphasis on safety and trustworthiness. Consequently, there has been significant interest in ensuring safety and trust for AI systems (Bostrom, 2014; Amodei et al., 2016; Russell, 2019; Jacovi et al., 2021; Tegmark and Omohundro, 2023).
In this work, we point out a fundamental tension between the requirements of an AI system being safe and trusted, but also matching or exceeding human capabilities, i.e. being an AGI system. There are several interpretations of safety, trust and AGI and our result does not preclude achieving these desiderata simultaneously under more relaxed interpretations that could still be useful in many practical applications. Therefore, to understand the limitations pointed out by our result, it is important to first understand our formalizations of these notions, and we will immediately proceed with defining these notions. We start by first defining an AI system for a given task.
@@
Definition 1.1 simply formalizes the notion of a system which solves instances of a task. In this paper, we will consider the tasks of program verification, planning and determining graph reachability (defined rigorously later). Note that we allow the system to abstain from providing an answer for some instance if it so determines, which could be important from the perspective of safety (Geifman and El-Yaniv, 2017). Now, we define the notion of safety.
Definition 1.2 (Safety).
We define a system to be safe if it does not make any false claims, i.e., for every instance it either answers the instance correctly or abstains from answering it.
As an example, in the context of verifying that a program has some specified property (such as always terminating), the system is safe if it does not classify a program as having the desired property if it does not have that particular property. Our definition allows the system to abstain from answering an instance if it is uncertain, but it requires the system to be correct whenever it outputs an answer. While less stringent definitions may suffice for some tasks, small probabilities of error may not be tolerable for mission-critical tasks, especially as system capabilities grow (Amodei et al., 2016; Tegmark and Omohundro, 2023). Next, we define trust as simply the assumption of safety.
@@
We define trust to be the assumption that the system is safe.
@@
To elaborate on the definition, if a system is trusted then it is scientifically accepted (or assumed) that the system is safe. As a remark, we note that our results are agnostic to whether trust in the system stems from theoretical proofs, empirical verification, or some combination of these, we only require that when deploying the system there is an assumption that it is safe. We also note that safety does not necessarily imply trust, or vice versa. Safety is an underlying property of the system being consistent and not making false claims. It is possible that some analysis of the system cannot identify this property or is incorrect, leading to a lack of trust or mistaken trust. For example, a system could actually be safe but not trusted because existing empirical or theoretical tools are insufficient to establish safety. Similarly, a system could actually be unsafe but still trusted by users, such as when the trust rests on empirical evidence which is incomplete, or on incorrect theoretical assumptions.
Finally, we need to formally define an AGI system in order to mathematically investigate its limitations. This is a challenge, since it is well-accepted that there is no well-accepted definition of AGI—or even of intelligence itself (Legg and Hutter, 2007; Legg et al., 2007). Nevertheless, we propose a formal definition, and argue why it captures important aspects of the goal of AGI.
Definition 1.4 (AGI).
@@
We define a system to be an Artificial General Intelligence (AGI) system if for every task instance such that a human has a provably correct solution for that instance, the system can also solve the instance with some non-zero probability. Similarly, the system is not an AGI system if there exists some task instance which can be easily and provably solved by a human, but the system can never solve the instance (for probabilistic systems, the probability of the system solving the instance is 0).
@@
Our definition draws on the common view that an AGI system for a task such as program verification should be at least as capable as a human on that task. In particular, if there are explicit task instances which can be provably solved by humans (for example, explicit programs which the humans can easily and provably certify as having the desired property) but cannot be solved by the system, then the system is not an AGI system as per our definition. We note that our definition bears some similarity to notions of superintelligence (Bostrom, 2014; Morris et al., 2024), and the reader can regard Definition 1.4 as a definition of superintelligence if they so prefer. For example Bostrom (2014) defines superintelligence as “any intellect that greatly exceeds the cognitive performance of humans in virtually all domains of interest”, and Morris et al. (2024) defines Level 5 AGI, which they term artificial superintelligence, as “outperforming 100% of humans” on a “wide range of non-physical tasks”. One distinction between these notions of superintelligence and our definition of AGI is that our definition does not require the AI system to necessarily outperform humans, but it does require the system to do at least as well as humans on all task instances.
@@
In our definition, when we say that a human has a provably correct solution, we mean that the human can provide a scientifically acceptable proof. In this paper whenever we make claims about humans being able to solve problems, we provide such proofs. To probe this point and Definition 1.4 further, we consider an analogy to chess — a domain for which we have had advanced AI systems for quite some time. Consider a future proposed AGI system, which is proficient at chess among other things. If there were explicit chess positions which most human chess players can solve provably without too much difficulty, but the proposed AGI system struggled on those positions, then the proposed system does not capture some aspects of human cognition, and hence is arguably not actually an AGI system. [^1] We note that many current advanced chess engines still struggle to evaluate certain positions which are relatively easy for human experts (Doggers, 2017; Zahavy et al., 2023). However, this is likely a result of the these engines being ‘narrow’ in terms of their approach and reasoning, and we believe that a proposed AGI or superintelligent system which is purported to excel on chess should have the ability to solve such instances. Similarly, in our paper we will demonstrate explicit instances of certain tasks for which we provide solutions with short, scientifically acceptable proofs which are also rather simple, but these instances cannot be solved by AI systems having certain properties.
@@
We now state our main result, that it is not possible for an AGI system to be both safe and trusted, as per our definitions of safety, trust and AGI. In other words, the notions of safety, trust and AGI are mutually incompatible — any system can have at most two of these three properties.
@@
If an AI system is safe and trusted, then it cannot be an AGI system. In particular, it is not an AGI system for the tasks of program verification, planning and determining graph reachability.
Theorem 1.5 points out a fundamental limitation of an AGI system: such a system cannot be both safe and trusted. Similarly, if there is some trusted AI system, then either that system is not actually safe, or it is not an AGI system. We prove this result in Section 3. While much of our proof technique mimics Gödel’s proof of his incompleteness theorems (Gödel, 1931) (and also Turing’s proof of the undecidability of the halting problem (Turing, 1937)), the argument we make is not in the context of axiomatic system and theorem proving but in the context of an AI system that needs to solve certain task instances of applications such as program verification or planning. Our proofs are self-contained in this context and do not require knowledge of formal axiomatic reasoning or logical rules of deduction. Thus rather than viewing the results as limitations of systems of logic, they should be viewed as limitations of AI systems.
We also consider a relaxation of safety which requires the AI system to be calibrated with respect to its predictions, as opposed to Definition 1.2 which requires the system to be always correct unless it abstains. We call this notion calibration-safety, and for the case of program verification calibration-safety requires that if the AI system outputs that some program terminates with some probability $p$, then that program should actually terminate with probability approximately $p$. In Section 4 we show a similar limitation as in Theorem 1.5 for AI systems which satisfy calibration-safety.
@@ 2 Related Work
In this section, we discuss some more related work on AGI, safety and trust in AI, and limitations of AI in the context of Gödel’s results.
Artificial General Intelligence.
Though not termed as “Artificial General Intelligence (AGI)” until more recently (Goertzel and Pennachin, 2007), the concept of machines which match or surpass the cognitive capabilities of humans dates back to the earliest days of AI (Turing, 1950; McCarthy et al., 1955; Minsky, 1961). Due to recent advances in foundation models such as large language models (Bommasani et al., 2021), there has been significant interest and capital investments in developing systems capable enough to be termed as an AGI both from the private sector and from governments (Maslej et al., 2025).
Safety and Trust in AI.
Safety concerns around advanced AI systems similarly date back to early days of AI (Turing, 1951; Wiener, 1950). With growing system capabilities, there has been significant recent focus on ensuring safety and trust to manage risks associated with AI systems (Future of Life Institute, 2024; Center for AI Safety, 2025). We refer the interested reader to several recent surveys and roadmaps for ensuring safety and trust in highly-capable, general purpose AI systems (Bengio et al., 2024; Chua et al., 2024; Chen et al., 2024; Bengio et al., 2025). It is also important to recognize that AI safety and trust encompass many facets beyond those considered in our definitions. For example, even formally specifying safety objectives can be challenging for complex tasks (Amodei et al., 2016), which introduces additional challenges to develop safe AI systems beyond those pointed out in our work.
@@
Fundamental limits on theorem proving and program verification were famously established by Gödel’s incompleteness theorems and Turing’s undecidability results. Gödel showed that in any sufficiently expressive formal system, there exist true statements (also called Gödel statements) that cannot be formally proven within the system (Gödel, 1931). Building on this, Turing proved that the Halting Problem—determining whether an arbitrary program halts on a given input—is undecidable (Turing, 1937), meaning no algorithm can solve it for all possible programs. These results imply that fully automatic verification of arbitrary program behavior, such as ensuring termination, is provably impossible in the general case. Our result uses similar ideas to draw a separation between the abilities of a safe, trusted AI system and humans.
@@
The goal of our work is distinct from that of Penrose and Lucas, and we do not aim to show a separation between any AI system and human reasoning. Instead, we prove a more restricted but rigorous result: that safe, trusted AI systems (under formal definitions of those terms) are necessarily unable to solve certain problems that humans can solve with provable correctness. The assumption of safety and trust is crucial (as will be evident from our proofs) — it allows humans to conclude the correctness of some outputs even when the AI system, by its own constraints, must abstain.
@@ 3 Technical Results
In this section, we discuss our main technical results regarding limitations of safe, trusted, AGI for program verification, planning, and graph reachability.
@@ 3.1 Program verification
procedure $\mathtt{G\ddot{o}del\_program}$   if $A(\mathtt{G\ddot{o}del\_program})==\text{`well-behaved'}$ then    while true do    end while   else    return 0   end if end procedure Claim (informal): If $A$ is safe then $\mathtt{G\ddot{o}del\_program}$ is well-behaved, but $A$ cannot output that $\mathtt{G\ddot{o}del\_program}$ is well-behaved. Proof sketch:
If $A$ outputs that $\mathtt{G\ddot{o}del\_program}$ is well-behaved, then the program enters an infinite loop. If $A$ is safe, this is a contradiction, hence $A$ cannot output $\mathtt{G\ddot{o}del\_program}$ is well-behaved. If $A$ does not output $\mathtt{G\ddot{o}del\_program}$ is well-behaved, then the program immediately terminates and hence is well-behaved.
Figure 1: Sketch of the basic argument for program verification, for the case when the AI system $A$ is well-behaved (i.e., always terminates) and deterministic. If $A$ is safe, it cannot determine if $\mathtt{G\ddot{o}del\_program}$ is well-behaved. However, since the condition of $A$ being safe is satisfied for a trusted system, it is possible to prove that $\mathtt{G\ddot{o}del\_program}$ is well-behaved for a trusted system. Therefore, a safe, trusted system cannot solve this instance, even though it is provably solvable.
@@
We define a program to be well-behaved if it terminates on every input (for randomized programs, the program terminates with probability 1). In the program verification problem, the system is given a program instance and it classifies the instance as being ‘well-behaved’, ‘not well-behaved’ or abstains from making a prediction (outputs ‘don’t know’). Safety for program verification requires that the system never outputs that a well-behaved program is not well-behaved, and vice versa. The system is trusted if we assume that the system is safe. Note that the system is not an AGI system if there is a well-behaved program which can be easily proven to be well-behaved by a human, but for which the program always abstains from making a prediction.
Our definition of program verification (Definition 3.1) and our results are for the property of the program halting. We believe it is possible to extend the result for verifying other semantic properties of programs — analogous to Rice’s theorem (Rice, 1953). We now state our result for program verification.
@@
If a system is safe and trusted, then it cannot be an AGI system for program verification.
@@
Note that $\mathtt{G\ddot{o}del\_program}$ involves running the program $P$ on the input pair $(P,T)$. $\mathtt{G\ddot{o}del\_program}$ checks that $(P,T)$ is a valid input type to the program $P$, and we can also regard the input $(P,T)$ as one input to $P$ that is a pair of entities: a program $P$ and a trace $T$. We now show that if $A$ is safe, then $\mathtt{G\ddot{o}del\_program}$ is well-behaved.
@@
If $A$ is safe, then $\mathtt{G\ddot{o}del\_program}$ is well-behaved.
@@
Now consider the case when the else if condition in step 4 is satisfied. This happens when the trace $T$ concludes that $P$ is well-behaved, and since the system $A$ is safe, then $P$ must be well-behaved in that case. Hence the execution in step 6 always terminates, and hence $\mathtt{G\ddot{o}del\_program}$ terminates for that input.
On the other hand, if the trace does not conclude that $P$ is well-behaved, then the program enters the else condition in lines 8 and 9, and immediately terminates. Therefore if $A$ is safe, then $\mathtt{G\ddot{o}del\_program}$ is well-behaved. ∎
We now note that the assumption of $A$ being safe is satisfied for a trusted system $A$. Therefore, the proof of Lemma 3.3 provides a short proof that $\mathtt{G\ddot{o}del\_program}$ is well-behaved for a safe and trusted system. In other words, if it is scientifically accepted than $A$ is safe, then the proof of Lemma 3.3 provides a scientifically acceptable proof that $\mathtt{G\ddot{o}del\_program}$ is well-behaved. Next, we show that the system $A$ cannot solve this instance.
@@
Therefore, if $A$ is safe and trusted, then it cannot be an AGI system. ∎

### v2

# Limitations on Accurate, Trusted, Human-level Reasoning

@@ Abstract
We identify a fundamental incompatibility between the goals of accuracy, trust, and human-level reasoning in artificial intelligence (AI) systems, for strict mathematical definitions of these notions. We define accuracy of a system as the property that it never makes any false claims when it has the ability to abstain from making a prediction on any input, and trust as the assumption that the system is accurate. We define human-level reasoning as the property of an AI system always matching or exceeding human capability. Our core finding is that—for our formal definitions of these notions—an accurate and trusted AI system cannot be a human-level reasoning system: for such an accurate, trusted system there are task instances which are easily and provably solvable by a human but not by the system. Our proofs draw parallels to Gödel’s incompleteness theorems and Turing’s proof of the undecidability of the halting problem, and can be regarded as interpretations of Gödel’s and Turing’s results. Key to our proof is the formalization of the notion of trust, which allows us to separate the intrinsic property of a system (being accurate) from its epistemic status (being trusted).
@@ 1 Introduction
Rapid advancements in artificial intelligence have intensified focus on achieving human-level reasoning across diverse tasks (Morris et al., 2024; Feng et al., 2024). AI systems capable of human-level reasoning have the potential for vast societal benefits through transformative impacts on nearly every aspect of society, including healthcare (Singhal et al., 2025), scientific research (Wang et al., 2023), education (Wang et al., 2024), sustainability (Rolnick et al., 2022), and economic growth (Chui et al., 2023). At the same time, development of such powerful systems necessitates a foundational emphasis on accuracy, reliability and trustworthiness. Consequently, there has been significant interest in ensuring accuracy, reliability and trust for AI systems (Bostrom, 2014; Amodei et al., 2016; Russell, 2019; Jacovi et al., 2021; Tegmark and Omohundro, 2023).
In this work, we point out a fundamental tension between the requirements of an AI system being accurate and trusted, but also matching or exceeding human reasoning capabilities, i.e. being a human-level reasoning system. There are several interpretations of accuracy, trust and human-level reasoning and our result does not preclude achieving these desiderata simultaneously under more relaxed interpretations that could still be useful in many practical applications. Therefore, to understand the limitations pointed out by our result, it is important to first understand our formalizations of these notions, and we will immediately proceed with defining these notions. We start by first defining an AI system for a given task.
@@
Definition 1.1 simply formalizes the notion of a system which solves instances of a task. In this paper, we will consider the tasks of program verification, planning and determining graph reachability (defined rigorously later). Note that we allow the system to abstain from providing an answer for some instance if it so determines, which could be important from the perspective of reliability and safety (Geifman and El-Yaniv, 2017). Also note that the definition does not require the AI system to necessarily provide a proof of its answer, the system only needs to provide an answer or abstain. Next, we define the notion of accuracy.
Definition 1.2 (Accuracy).
We define a system to be accurate if it does not make any false claims, i.e., for every instance it either answers the instance correctly or abstains from answering it. For a randomized system, if the system answers the instance with some non-zero probability instead of always abstaining, then its answer must be correct.
As an example, in the context of verifying that a program has some specified property (such as always terminating), the system is accurate if it does not classify a program as having the desired property if it does not have that particular property. Our definition allows the system to abstain from answering an instance if it is uncertain, but it requires the system to be correct whenever it outputs an answer. Note that our notion of accuracy is closely related to notions of factuality, reliability and, in some situations, even safety. Often abstention is preferred over a confident but incorrect response as the consequences of an incorrect response may be severe. Even small probabilities of error may not be tolerable for mission-critical tasks, especially as system capabilities grow (Amodei et al., 2016; Tegmark and Omohundro, 2023).
Next, we define trust. Trust is defined with respect to some set of individuals $H$, and is the assumption of accuracy.
@@
We define a system to be trusted by $H$ if for every $\in H, hh$ has the assumption that the system is accurate.
@@
Therefore, if a system is trusted by $H$ then it is assumed by any individual in $H$ that the system is accurate. For brevity, we will sometimes refer to a system as being “trusted” without explicitly referencing the set $H$, if there exists some set of individuals $H$ such that the system is trusted by $H$. As a remark, we note that our results are agnostic to whether trust in the system stems from theoretical proofs, empirical verification, or some combination of these, we only require that when using the system there is an assumption that it is accurate. We also note that accuracy does not necessarily imply trust, or vice versa. Accuracy is an underlying property of the system being consistent and not making false claims. It is possible that some analysis of the system cannot identify this property or is incorrect, leading to a lack of trust or mistaken trust. For example, a system could actually be accurate but not trusted because existing empirical or theoretical tools are insufficient to establish accuracy. Similarly, a system could actually be inaccurate but still trusted by users, such as when the trust rests on empirical evidence which is incomplete, or on incorrect theoretical assumptions.
Finally, we need to formally define a human-level reasoning system in order to mathematically investigate its limitations.
Definition 1.4 (Human-level reasoning).
@@
We define a system to be a human-level reasoning system if for every task instance such that a human has a provably correct solution for that instance, the system can also solve the instance with some non-zero probability. Similarly, the system is not a human-level reasoning system if there exists some task instance which can be easily and provably solved by a human, but the system can never solve the instance (for probabilistic systems, the probability of the system solving the instance is 0).
@@
Our definition draws on the common view that a human-level reasoning system for a task such as program verification should be at least as capable as a human on that task. In particular, if there are explicit task instances which can be provably solved by humans (for example, explicit programs which the humans can easily and provably certify as having the desired property) but cannot be solved by the system, then the system is not a human-level reasoning system as per our definition.
Our definition draws on some similar notions of artificial general intelligence (AGI). It is well-accepted that there is no well-accepted definition of AGI—or even of intelligence itself (Legg and Hutter, 2007; Legg et al., 2007)—but the term AGI is usually understood to mean that the AI system should be at least as capable as humans across diverse tasks. The term superintelligence is also used in the context of AGI (Bostrom, 2014; Morris et al., 2024). For example Bostrom (2014) defines superintelligence as “any intellect that greatly exceeds the cognitive performance of humans in virtually all domains of interest”, and Morris et al. (2024) defines Level 5 AGI, which they term artificial superintelligence, as “outperforming 100% of humans” on a “wide range of non-physical tasks”. One distinction between these notions of superintelligence and our definition of human-level reasoning is that our definition does not require the AI system to necessarily outperform humans, but it does require the system to do at least as well as humans on all task instances.
@@
In our definition, when we say that a human has a provably correct solution, we mean that the human can provide a proof. In this paper whenever we make claims about humans being able to solve problems, we provide such proofs. To probe this point and Definition 1.4 further, we consider an analogy to chess — a domain for which we have had advanced AI systems for quite some time. Consider a future proposed human-level reasoning system, which is proficient at chess among other things. If there were explicit chess positions which most human chess players can solve provably without too much difficulty, but the proposed human-level reasoning system struggled on those positions, then the proposed system does not capture some aspects of human cognition, and hence is arguably not actually a human-level reasoning system. [^1] We note that many current advanced chess engines still struggle to evaluate certain positions which are relatively easy for human experts (Doggers, 2017; Zahavy et al., 2023). However, this is likely a result of the these engines being ‘narrow’ in terms of their approach and reasoning, and we believe that a proposed human-level reasoning or superintelligent system which is purported to excel on chess should have the ability to solve such instances. Similarly, in our paper we will demonstrate explicit instances of certain tasks for which we provide solutions with short proofs which are also rather simple, but these instances cannot be solved by AI systems having certain properties.
@@
We now state our main result, that it is not possible for a human-level reasoning system to be both accurate and trusted, as per our definitions of accuracy, trust and human-level reasoning. In other words, the notions of accuracy, trust and human-level reasoning are mutually incompatible — any system can have at most two of these three properties.
@@
If an AI system is accurate and trusted by some set of individuals $H$, then it cannot be a human-level reasoning system. In particular, it is not a human-level reasoning system for the tasks of program verification, planning and determining graph reachability.
Theorem 1.5 points out a fundamental limitation of a human-level reasoning system: such a system cannot be both accurate and trusted. Similarly, if there is some trusted AI system, then either that system is not actually accurate, or it is not a human-level reasoning system. We prove this result in Section 3. While much of our proof technique mimics Gödel’s proof of his incompleteness theorems (Gödel, 1931) (and also Turing’s proof of the undecidability of the halting problem (Turing, 1937)), the argument we make is not in the context of axiomatic system and theorem proving but in the context of an AI system that needs to solve certain task instances of applications such as program verification or planning. Our proofs are self-contained in this context and do not require knowledge of formal axiomatic reasoning or logical rules of deduction. Thus rather than viewing the results as limitations of systems of logic, they should be viewed as limitations of AI systems.
We also consider a relaxation of accuracy which requires the AI system to be calibrated with respect to its predictions, as opposed to Definition 1.2 which requires the system to be always correct unless it abstains. In the context of program verification, calibration requires that if the AI system outputs that some program terminates with probability $p$, then that program should actually terminate with probability approximately $p$. In Section 5 we show a similar limitation as in Theorem 1.5 for AI systems which are calibrated.
@@ 2 Related Work
In this section, we discuss some more related work on human-level reasoning, accuracy and trust in AI, and limitations of AI in the context of Gödel’s results.
Human-level Reasoning Systems.
Though not termed as “Artificial General Intelligence (AGI)” until more recently (Goertzel and Pennachin, 2007), the concept of machines which match or surpass the cognitive capabilities of humans dates back to the earliest days of AI (Turing, 1950; McCarthy et al., 1955; Minsky, 1961). Due to recent advances in foundation models such as large language models (Bommasani et al., 2021), there has been significant interest and capital investments in developing systems capable of human-level reasoning both from the private sector and from governments (Maslej et al., 2025).
Accuracy, Reliability and Trust in AI.
Concerns around risks associated with advanced AI systems similarly date back to early days of AI (Turing, 1951; Wiener, 1950). With growing system capabilities, there has been significant recent focus on ensuring safety and trust in the context of AI systems (Future of Life Institute, 2024; Center for AI Safety, 2025). We refer the interested reader to several recent surveys and roadmaps for ensuring safety and trust in highly-capable, general purpose AI systems (Bengio et al., 2024; Chua et al., 2024; Chen et al., 2024; Bengio et al., 2025). It is also important to recognize that AI safety and trust encompass many facets beyond those considered in our definitions. For example, even formally specifying safety objectives can be challenging for complex tasks (Amodei et al., 2016), which introduces additional challenges to develop safe AI systems beyond those pointed out in our work.
@@
Fundamental limits on theorem proving and program verification were famously established by Gödel’s incompleteness theorems and Turing’s undecidability results. Gödel showed that in any sufficiently expressive formal system, there exist true statements (also called Gödel statements) that cannot be formally proven within the system (Gödel, 1931). Building on this, Turing proved that the Halting Problem—determining whether an arbitrary program halts on a given input—is undecidable (Turing, 1937), meaning no algorithm can solve it for all possible programs. These results imply that fully automatic verification of arbitrary program behavior, such as ensuring termination, is provably impossible in the general case. Our result uses similar ideas to draw a separation between the abilities of an accurate, trusted AI system and humans.
@@
The goal of our work is distinct from that of Penrose and Lucas, and we do not aim to show a separation between any AI system and human reasoning. Instead, we prove a more restricted but rigorous result: that accurate, trusted AI systems (under formal definitions of those terms) are necessarily unable to solve certain problems that humans can solve with provable correctness. The assumption of accuracy and trust is crucial (as will be evident from our proofs) — it allows humans to conclude the correctness of some outputs even when the AI system, by its own constraints, must abstain.
@@ 3 Technical Results
In this section, we discuss our main technical results regarding limitations of accurate, trusted, human-level reasoning for program verification, planning, and graph reachability.
@@ 3.1 Program verification
procedure $\mathtt{G\ddot{o}del\_program}$   if $A(\mathtt{G\ddot{o}del\_program})==\text{`well-behaved'}$ then    while true do    end while   else    return 0   end if end procedure Claim (informal): If $A$ is accurate then $\mathtt{G\ddot{o}del\_program}$ is well-behaved, but $A$ cannot output that $\mathtt{G\ddot{o}del\_program}$ is well-behaved. Proof sketch:
If $A$ outputs that $\mathtt{G\ddot{o}del\_program}$ is well-behaved, then the program enters an infinite loop. If $A$ is accurate, this is a contradiction, hence $A$ cannot output $\mathtt{G\ddot{o}del\_program}$ is well-behaved. If $A$ does not output $\mathtt{G\ddot{o}del\_program}$ is well-behaved, then the program immediately terminates and hence is well-behaved.
Figure 1: Sketch of the basic argument for program verification, for the case when the AI system $A$ is well-behaved (i.e., always terminates) and deterministic. If $A$ is accurate, it cannot determine if $\mathtt{G\ddot{o}del\_program}$ is well-behaved. Now if $A$ is trusted by the set $H$ then individuals in $H$ assume that $A$ is accurate, and hence the condition ‘If AA is accurate’ is satisfied. Therefore if $A$ is accurate and trusted by $H$, then individuals in $H$ can prove that $\mathtt{G\ddot{o}del\_program}$ is well-behaved by the proof sketched above. Therefore, an accurate, trusted system cannot solve this instance, even though it is provably solvable by individuals in $H$.
@@
We define a program to be well-behaved if it terminates on every input (for randomized programs, the program terminates with probability 1). In the program verification problem, the system is given a program instance and it classifies the instance as being ‘well-behaved’, ‘not well-behaved’ or abstains from making a prediction (outputs ‘don’t know’). Accuracy for program verification requires that the system never outputs that a well-behaved program is not well-behaved, and vice versa. The system is trusted if we assume that the system is accurate. Note that the system is not a human-level reasoning system if there is a well-behaved program which can be easily proven to be well-behaved by a human, but for which the program always abstains from making a prediction.
Our definition of program verification (Definition 3.1) and our results are for the property of the program halting. In Section 4, we extend the result to a general class of properties of programs, including those relevant from the perspective of safety. We now state our result for program verification.
@@
If a system is accurate and trusted, then it cannot be a human-level reasoning system for program verification.
@@
Note that $\mathtt{G\ddot{o}del\_program}$ involves running the program $P$ on the input pair $(P,T)$. $\mathtt{G\ddot{o}del\_program}$ checks that $(P,T)$ is a valid input type to the program $P$, and we can also regard the input $(P,T)$ as one input to $P$ that is a pair of entities: a program $P$ and a trace $T$. We now show that if $A$ is accurate, then $\mathtt{G\ddot{o}del\_program}$ is well-behaved.
@@
If $A$ is accurate, then $\mathtt{G\ddot{o}del\_program}$ is well-behaved.
@@
Now consider the case when the else if condition in step 4 is satisfied. This happens when the trace $T$ concludes that $P$ is well-behaved, and since the system $A$ is accurate, then $P$ must be well-behaved in that case. Hence the execution in step 6 always terminates, and hence $\mathtt{G\ddot{o}del\_program}$ terminates for that input.
On the other hand, if the trace does not conclude that $P$ is well-behaved, then the program enters the else condition in lines 8 and 9, and immediately terminates. Therefore if $A$ is accurate, then $\mathtt{G\ddot{o}del\_program}$ is well-behaved. ∎
Note that if $A$ is trusted by some set $H$ then individuals in $H$ assume that $A$ is accurate. Therefore if $A$ is accurate and trusted by $H$ then individuals in $H$ can prove that $\mathtt{G\ddot{o}del\_program}$ is well-behaved, indeed the proof of Lemma 3.3 provides a short proof of this since the condition of $A$ being accurate is satisfied due to trust. Next, we show that the system $A$ cannot solve this instance.
@@
Therefore, if $A$ is accurate and trusted, then it cannot be a human-level reasoning system. Note that the proof illustrates a sharp tension between trust and human-level reasoning for an accurate system. As long as some set of individuals trust an accurate system $A$, they can provably solve a task instance which cannot be solved by $A$. If there is a large set of individuals who trust the accurate system $A$, then a large set of individuals can solve a task instance which cannot be solved by $A$. ∎
A few notes about the results and the setting are in order.
  • The result is agnostic to the type of system $A$ (e.g. $A$ being a large language model or otherwise), it only requires $A$ to be programmatically defined so that $\mathtt{G\ddot{o}del\_program}$ is a valid program. We also note that the result in fact holds even if $A$ is a non-deterministic Turing machine, since it only concerns traces of the system $A$.
  • The result also holds for randomized systems. It establishes that an accurate, randomized system $A$ cannot output ‘well-behaved’ on $\mathtt{G\ddot{o}del\_program}$ with any probability greater than 0 (since otherwise there is a trace which concludes ‘well-behaved’, in which case the output of $\mathtt{G\ddot{o}del\_program}$ is different from its own output).
We discuss some further, higher-level remarks about our result in the discussion in Section 6.